OpenStudy (anonymous):

HELPPPPP!!!!! Use the rules of exponents to evaluate or simplify. Write without negative exponents. y^-5/y^-2

OpenStudy (campbell_st):

well you can use the exponent rule for division 1st, which is subtract the powers \[\frac{x^a}{x^b} = x^{a - b}\] then use the exponent rule for fractions... \[x^{-a} = \frac{1}{x^a}\] hope this helps.

OpenStudy (anonymous):

flip it meanning ... y^2 /Y^5 when you flip it the negitives go away

OpenStudy (anonymous):